Solve for all values of X.
(2x over x+3)+ (3 over x-3) = (8 over x squared - 9)

Answer by fanism(39)

(2x over x+3)+ (3 over x-3) = (8 over x squared - 9)
[2x(x-3) + 3(x+3)]/(x^2 -9 )= 8 / (x^2-9)
2x^2-6x+3x+9=8
2x^2-3x+9=8
2x^2-3x+1=8
(2x-1)(x-1)=0
solve x yourself